A motor scooter travels east at a speed of 13 m/s. The driver then reverses direction and heads west at 17 m/s.  What was the change in velocity of the scooter?

Grade:12

2 Answers

suryakanth AskiitiansExpert-IITB
13 years ago

Dear sachin,

the velocity of scooter in vector form is

V1 = 13i

Vfinal = -17 m/s

change in velocity = Vf - V intial

  = -17-13 = -30 m/s

Hence change in velocity = -30 m/s

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Suryakanth –IITB

rajiv kumar
13 years ago

negative of 20 m/s.